Assume that both Italy and France can produce cheese and wine. Italian workers can produce either 200 cases of wine or 600 pounds of cheese in one month and French workers can produce 250 cases of wine or 500 pounds of cheese in one month. Assume constant opportunity costs.

  1. Graph the PPF/PPC for each country on a single graph
  2. Which country has the absolute advantage in wine? ___________
  3. Which country has the absolute advantage in cheese? __________
  4. What is the opportunity cost of a case of wine in Italy? _________
  5. What is the opportunity cost of a case of wine in France? ________
  6. What is the opportunity cost of a pound of cheese in Italy? ______
  7. What is the opportunity cost of a pound of cheese in France? ______
  8. Which country has the comparative advantage in wine? __________
  9. Which country has the comparative advantage in cheese? __________
  10. On your PPF above, show the JOINT PPF
Answer #1
Wine (cases) Cheese (Pounds)
Italy 200 600
France 250 500

(a) By plotting the given PPF ,we get the graph as shown below:

A Cheese Pounds وما tooo be you Joint RRE 0 100 200 320 los soo Wine (lases)

(b) France has an absolute advantage in producing wine because it can produce more wine with the given resources than Italy.

(c) Italy has an absolute advantage in producing cheese because it can produce more cheese with the given resources than France.

(d) The opportunity cost of a case of wine in Italy = 600/200= 3 pounds of cheese.

(e) The opportunity cost of a case of wine in France = 500/250= 2 pounds of cheese.

(f) The opportunity cost of a pound of cheese in Italy = 200/600 = 0.33 cases of wine.

(g) The opportunity cost of a pound of cheese in France = 250/500= 0.5 cases of wine.

(h) France has a comparative advantage in wine because France has a lower opportunity cost for producing wine.

(i) Italy has a comparative advantage in cheese because Italy has a lower opportunity cost for producing cheese.

(j) Joint PP points are:

Wine = 450 cases

Cheese = 1100 pounds

It is shown in the above graph.